mathslover (mathslover):

The volume V of water passing any point of a uniform tube during t seconds is related to the cross-sectional area A of the tube and velocity u of water by the relation \[V \propto A^{\alpha} u^ {\beta} t^ {\gamma}\] Which one of the following will be true ?

OpenStudy (unklerhaukus):

\[[\textbf L]^3\propto [\mathbf L^2]^\alpha\cdot\frac{[\mathbf L]^\beta}{[\mathbf T]^\beta}\cdot[\mathbf T]^\gamma\]

mathslover (mathslover):

Yep Unkle. I then continued the process getting : \[2\alpha + \beta = 3 \] and \[-\beta + \gamma = 0 \]

mathslover (mathslover):

I think the situation or condition is satisfied at \(\alpha = \beta = \gamma\) but if we talk about the general equation or form, it is : \(\alpha \ne \beta = \gamma\)

mathslover (mathslover):

\(\large{\alpha \ne \beta = \gamma}\) is correct for each and every real value. But \(\alpha = \beta = \gamma\) satisfies only if all are equal to 1.

OpenStudy (unklerhaukus):

why do you say \(α≠β\)?

mathslover (mathslover):

\(\alpha \ne \beta\) because \(\alpha = \cfrac{3-\beta}{2}\) .

OpenStudy (unklerhaukus):

i dont understand that bit.

mathslover (mathslover):